# Provable In-Context Learning of Nonlinear Regression with Transformers

### v1

@@ Abstract
The transformer architecture, which processes sequences of input tokens to produce outputs for query tokens, has revolutionized numerous areas of machine learning. A defining feature of transformers is their ability to perform previously unseen tasks using task-specific prompts without updating parameters—a phenomenon known as in-context learning (ICL). Recent research has actively explored the training dynamics behind ICL, with much of the focus on relatively simple tasks such as linear regression and binary classification. To advance the theoretical understanding of ICL, this paper investigates more complex nonlinear regression tasks, aiming to uncover how transformers acquire in-context learning capabilities in these settings. We analyze the stage-wise dynamics of attention during training: attention scores between a query token and its target features grow rapidly in the early phase, then gradually converge to one, while attention to irrelevant features decays more slowly and exhibits oscillatory behavior. Our analysis introduces new proof techniques that explicitly characterize how the nature of general non-degenerate $L$-Lipschitz task functions affects attention weights. Specifically, we identify that the Lipschitz constant $L$ of nonlinear function classes as a key factor governing the convergence dynamics of transformers in ICL. Leveraging these insights, for two distinct regimes depending on whether $L$ is below or above a threshold, we derive different time bounds to guarantee near-zero prediction error. Notably, despite the convergence time depending on the underlying task functions, we prove that query tokens consistently attend to prompt tokens with highly relevant features at convergence, demonstrating the ICL capability of transformers for unseen functions.
@@ 1 Introduction
ICL has attracted growing interest in recent years, with numerous empirical studies aiming to understand the conditions under which transformers succeed or fail at in-context generalization [10, 11, 12, 13, 14, 15, 16]. Notably, [11] offered preliminary theoretical evidence that transformers trained on specific function classes (e.g., linear functions) can accurately infer the function of a query sample based on prompts of a sequence of variable and their corresponding function value pairs. This highlights the surprising ability of transformers to "learn" within their forward pass, mimicking classical function approximation.
Building on this foundation, a few works have provided theoretical understandings of ICL by characterizing the dynamics of transformer models with one attention layer during the training process [17, 18, 19, 20, 21, 22, 23]. For instance, [18] investigates convergence dynamics under gradient flow, but focuses solely on a single linear attention layer applied to linear regression. [19] extends this line of analysis to softmax attention, revealing how attention weights evolve during training. More recent studies consider nonlinear function classes, such as binary classification and low-degree polynomial regression [20, 22, 23], and characterize how transformers adapt to these tasks during training.
Despite recent advances, most theoretical studies of ICL remain limited to simplified function classes, such as linear or low-complexity mappings. However, understanding how transformers learn in practice demands moving beyond these idealized settings—insights derived from such simplified classes often fail to generalize to nonlinear functions commonly encountered in real-world tasks. While extending the analysis to more general function classes is crucial, it presents significant technical challenges due to the inherent complexity of nonlinear functions even for one-layer transformers.
In this work, we take a step toward closing this gap by developing a theoretical framework for ICL with more general nonlinear regression functions. We investigate two fundamental questions: (1) What salient geometric properties of the target function govern the convergence behavior of transformer-based learning? and (2) Despite the nonlinearity and generality of these functions, can the transformer still learn effectively in context and achieve globally small prediction error? We answer both questions by analyzing the training dynamics of transformers under gradient descent.
We summarize our key contributions as follows.
@@
    Broad Class of Nonlinear Functions and More General Feature Sets: Our analysis generalizes previous studies in two folds. (i) In contrast to prior works that focus on linear mappings [18, 19], binary classification [20], or low-degree polynomials [23], our analysis applies to much broader non-degenerate $L$-Lipschitz task functions without assuming low complexity. This class is satisfied by almost any nontrivial $L$-Lipschitz function (e.g., all non-constant linear functions) and excludes only degenerate cases. (ii) Our results also apply to more general feature embeddings without the restrictive orthonormality assumption used in prior work [19, 20, 25, 24].
@@
    Phase Transition of Training Dynamics between Flat and Sharp Curvature Regimes: We identify the Lipschitz constant $L$ of nonlinear function classes as a key factor governing the convergence dynamics of transformers in ICL. An interesting phase transition emerges between the flat and sharp curvature regimes, depending on whether $L$ is below or above a threshold of order $\Theta\left(\frac{1}{\Delta\delta}\right)$, with each regime exhibiting distinct training behavior. In the flat curvature regime, $L$ is below the threshold and induces smaller gradients, which allow for larger step sizes to convergence. In contrast, in the sharp curvature regime, $L$ is above the threshold and incurs large gradients, which necessitate smaller step sizes to converge. Comparing the two, the flat regime may yield faster convergence when a high level of accuracy is required. However, if $L$ is significantly larger relative to the desired accuracy, the increased curvature enhances feature distinguishability, enabling faster training in the sharp regime.
@@
    Convergence Guarantee for ICL with Nonlinear Functions: We provide the formal convergence guarantees for a one-layer softmax attention model trained via gradient descent to perform ICL over nonlinear regression functions determined by the Lipschitz constant $L$. For both flat and sharp $L$-regimes, we show that gradient descent converges to near-zero training loss in polynomial time. We further characterize the distinct convergence rates in the two regimes by analyzing a two-phase training dynamic: in the early phase, attention scores between query tokens and their target features grow rapidly; in the later phase, these scores gradually converge to one, while attention to irrelevant features decays more slowly with oscillations. Further, we prove that query tokens consistently attend to prompt tokens with highly relevant features at convergence, and hence the model can correctly predict the values for functions that have not been seen during the training, demonstrating the ICL capability.
@@
    Novel Analysis Techniques: Our analysis develops new proof techniques that explicitly characterize how the curvature of nonlinear task functions influences the evolution of attention weights during training. In particular, we show that the flat and sharp curvature regimes of the parameter $L$ lead to distinct gradient magnitudes, which in turn drive different convergence rates and shape the overall training dynamics. The impact of function curvature on the magnitude and stability of attention updates is, to our knowledge, not addressed in previous ICL theory literature.
@@ 2 Related Works
The Bayesian inference perspective, pioneered by [10] and further developed by [26, 27, 28], establishes a theoretical framework linking prompting strategies to probabilistic reasoning. This line of work interprets ICL as a form of implicit Bayesian model averaging, where transformers effectively perform approximated inference conditioned on the provided context. Another line of ICL work focuses on Markov chains to study the behavior of induction heads for transformers, which are designed to copy or compare tokens that follow previous occurrences ([24, 29, 30, 31, 32, 33, 34, 35]). The central focus is to understand how transformers recover latent sequence structures ([24]) and transition rules ([32, 33, 34, 35]), using token-level recurrence and dynamics. For example, [35] analyzes next-token prediction in first-order Markov chains, where the transformer is trained to learn transition dynamics in token sequences.
@@
In contrast, the function learning perspective initiated by [11] demonstrated transformers’ remarkable ability to learn and interpolate simple function classes (particularly linear models) directly from context examples. This work sparked significant interest in understanding the mechanistic underpinnings of ICL, leading to important discoveries by [12] and [36], who revealed deep connections between attention mechanisms and gradient-based optimization dynamics. Recent advances have substantially expanded our understanding of ICL mechanisms: [37] provided a rigorous analysis of linear regression tasks, showing that trained transformers can implement both ridge regression gradient descent and exact least-squares solutions. [13] characterized ICL as a two-stage process involving context encoding and task execution, explaining its strong empirical performance. [38] established comprehensive theoretical results encompassing expressive power, prediction capabilities, and sample complexity, while proposing general mechanisms for algorithmic selection.
@@
Theoretical analysis of ICL learning dynamics. Recent theoretical work has made significant progress in understanding the learning dynamics of in-context learning (ICL) in transformers, though important limitations remain [19, 20, 21, 22, 23, 39, 30, 40, 41, 16]. The foundational work by [19] established the first rigorous analysis of training dynamics for softmax attention in ICL settings, focusing on a single-head attention layer learning linear regression tasks. Their key theoretical result demonstrates that prompt tokens with features identical to the query token develop dominating attention weights during training. However, this analysis relies critically on strong assumptions about pairwise orthonormality of feature vectors, limiting its applicability to more general settings.
@@
Subsequent work extended these results to classification tasks [20], dual-head settings [40], and structured or hierarchical functions [22, 23]. Some studies explored the implicit bias of gradient descent and attention-based generalization [21, 39], while others examined information-theoretic and dynamical aspects of ICL [30, 41]. Despite these advances, most analyses still rely on restrictive assumptions, such as orthogonal features, fixed positional encoding, or carefully structured input distributions, limiting their ability to explain ICL under general tasks and practical learning conditions. While [42] interprets transformers as implementing functional gradient descent for nonlinear regression, their analysis primarily focuses on the representational capacity and functional viewpoint of the learned predictor, without analyzing the ICL learning dynamics.
@@
Notations. In this paper, for a vector $\bm{v}$, we let $\|\bm{v}\|_{2}$ denote its $\ell$-222 norm. For some positive constant $C_{1}$ and $C_{2}$, we define $x=\Omega(y)$ if $x>C_{2}|y|$, $x=\Theta(y)$ if $C_{1}|y|<x<C_{2}|y|$, and $x=\mathcal{O}(y)$ if $x<c_{1}|y|$. We also denote by $x=o(y)$ if $x/y\rightarrow 0$. We use $\mathrm{poly}(C)$ to denote large constant degree polynomials of $C$. For a matrix $A$, we use $A_{i}$ to denote the $i$-th column of $A$, and $A_{i:j}$ to represent the collection of columns from the $i$-th to the $j$-th column (inclusive).
@@ 3.1 In-Context Learning Problem Setup
We consider a standard in-context learning (ICL) framework commonly used in prior studies such as [11, 19, 22]. The objective is to train a transformer model that can perform ICL over a designated class of functions $\mathcal{F}$, where each function $f\in\mathcal{F}$ corresponds to one task context. Here, we focus on nonlinear function classes further elaborated below.
Given each task (i.e., given one function $f$ randomly sampled from $\mathcal{F}$), a prompt with a sequence of $N$ input-response pairs $(x_{i},y_{i})$ as well as a query input $x_{\mathrm{query}}$ are sampled, where $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, and $y_{i}=f(x_{i})$. Let the input matrix $X=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}\\ \end{pmatrix}\in\mathbb{R}^{d\times N}$ and the response vector $\mathbf{y}=\begin{pmatrix}y_{1}&y_{2}&\cdots&y_{N}\end{pmatrix}\in\mathbb{R}^{1\times N}$. We adopt the following standard prompt embedding ([11, 19, 22]):
$P=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathrm{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0\end{pmatrix}=\begin{pmatrix}X&x_{\mathrm{query}}\\ Y&0\end{pmatrix}\in\mathbb{R}^{(d+1)\times(N+1)}.$ (1)
@@
=${\mathcal{F}=\left\{f:\;\begin{array}[]{l}|f(x)-f(x^{\prime})|\leq L\|x-x^{\prime}\|,\quad\forall\|x-x^{\prime}\|=\Theta(\delta_{0}),\\[6.99997pt] \forall v_{k}\in\mathbb{V},\exists\,v_{k^{\prime}}\in\mathbb{V},\;k^{\prime}\neq k,\;\text{such that }|f(v_{k})-f(v_{k^{\prime}})|=\Theta(L)\cdot\|v_{k}-v_{k^{\prime}}\|\end{array}\right\},}$ (4)
where $L>0$ and $\delta_{0}=\mathcal{O}(1)$. In particular, the functions in the class are said to satisfy non-degenerate $L$-Lipschitz condition, which imposes two natural requirements on the function class. First, the standard global $L$-Lipschitz condition ensures that the function $f$ does not change too rapidly, which is common in the literature and includes a wide class of linear and nonlinear functions. Second, the separation requirement guarantees that for any feature $v_{k}$ in the set $\mathbb{V}$, there exists at least one other feature $v_{k^{\prime}}$ such that the function difference between them achieves the order of variation defined by the Lipschitz constant. This ensures sufficient distinguishability of features by the function class for guaranteed learnability.
This mild separation assumption is satisfied by almost any nontrivial $L$-Lipschitz function (e.g., all non-constant linear functions) and excludes only degenerate cases. Together, these two conditions ensure that the function class is sufficiently rich for ICL while avoiding unlearnable or trivial scenarios.
For each prompt $P$, the task-specific function $f(x)$ is independently drawn based on a task distribution $\mathcal{D}_{f}$, as long as $f(x)$ satisfies the property for the same $L$ and $\delta_{0}$ in (4).
Feature Embeddings. Let $\mathbb{V}:=\big{\{}v_{k}\in\mathbb{R}^{d}\big{|}k=1,\cdots,K\big{\}}$ be the feature embeddings of tokens. For any $k\neq k^{\prime}$, we assume a separation of $\|v_{k}-v_{k^{\prime}}\|=\Theta(\Delta)$, where $\Delta=\Theta(1)$. Each data sample $x$ is modeled as a noisy perturbation of one of the vectors in $\mathbb{V}$. Such a condition can be satisfied by various feature learning techniques to avoid feature collapse, e.g., disentangled representation learning [43, 44, 45]. We note that such a condition substantially generalizes the orthogonality assumption taken by the previous study [19, 20, 25, 24], which assumes that all feature embeddings are orthogonal to each other.
The prompt is sampled as follows. For a randomly chosen $v_{k}$, we assume $x$ satisfies $\|x-v_{k}\|=O(\epsilon_{x})$ with probability $p_{k}$, where $\epsilon_{x}=o(1)$ and $p_{k}=\Theta(\frac{1}{K})$. For analytical simplicity, we assume $x=v_{k}$ whenever this proximity condition holds.

### v2

@@ Abstract
The transformer architecture has revolutionized machine learning by processing input sequences into outputs. A defining feature is in-context learning (ICL)—the ability to perform unseen tasks from prompts without updating model parameters. Early theoretical work focused on linear tasks, and recent studies have begun exploring nonlinear functions. Yet a rigorous analysis of the training dynamics—how transformers learn such complex tasks—remains elusive. This paper presents the first formal analysis of ICL training dynamics for a broad class of nonlinear regression functions. We analyze the stage-wise dynamics of attention during training: attention scores between a query token and its target features rise rapidly at first, then gradually converge to one, while attention to irrelevant features decays more slowly and can oscillate. Our analysis explicitly characterizes how general non-degenerate $L$-Lipschitz task functions shape attention weights, identifying the Lipschitz constant $L$ as the key factor governing the convergence dynamics. Leveraging these insights, for two distinct regimes depending on whether $L$ is below or above a threshold, we derive different time bounds to guarantee near-zero prediction error. Despite convergence time depending on the task, we prove query tokens ultimately focus on highly relevant prompt tokens, demonstrating transformers’ robust ICL capability for unseen functions.
@@ 1 Introduction
ICL has attracted growing interest, with numerous empirical studies examining when transformers succeed or fail at in-context generalization [10, 11, 12, 13, 14, 15, 16]. Notably, [11] provided preliminary theoretical evidence that transformers trained on specific function classes (e.g., linear) can accurately infer a query function value from prompts containing variable–function pairs, highlighting transformers’ surprising ability to “learn” within their forward pass and mimic classical function approximation.
Building on this foundation, subsequent works have provided theoretical understandings of ICL by characterizing the training dynamics of single-layer attention transformers [17, 18, 19, 21, 22]. For instance, [19] consider softmax attention to analyze how attention weights evolve during training linear regression problems. More recent studies have theoretically shown that transformers can learn specific nonlinear function classes in-context, such as binary classification, low-degree polynomial regression, and Gaussian single-index models [20, 22, 27, 23]. However, these studies do not provide a full theoretical picture of how the step-by-step learning process is governed by the task itself. To date, a formal characterization of the pre-training dynamics for general nonlinear ICL has been a key open problem.
In this work, we take a step toward understanding the learning dynamics for ICL on a broad class of nonlinear regression functions. We address two fundamental questions: (1) Which geometric properties of the target function govern the convergence behavior of transformer-based ICL? and (2) Despite nonlinearity and generality, how can a transformer learn in context to achieve low prediction error? We answer both by analyzing transformer training under gradient descent. Our main contributions are summarized below.
@@
    Broad Class of Nonlinear Functions and Flexible Feature Sets: Our analysis generalizes previous studies in two ways. (i) Unlike prior theoretical works that focus on linear mappings [18, 19], binary classification [20], or low-degree polynomials [23], we characterize learning dynamics for a much broader family of non-degenerate $L$-Lipschitz task functions without assuming low complexity. This class is satisfied by almost any nontrivial $L$-Lipschitz function (e.g., scaled cosine, piecewise-polynomial, and small neural-network functions) and excludes only degenerate cases. (ii) Our results also hold for general feature embeddings without the restrictive orthonormality assumptions in prior work [19, 20, 25, 24].
@@
    Phase Transition of Training Dynamics between Flat and Sharp Curvature Regimes: We discover a phase transition in training dynamics governed by the Lipschitz constant $L$. When $L$ is below a threshold of order $\Theta\left(\frac{1}{\Delta\delta}\right)$, the flat curvature regime yields smaller gradients and permits larger step sizes to converge. When $L$ exceeds the threshold, the sharp-curvature regime produces larger gradients requiring smaller steps. The two regimes exhibit distinct convergence behaviors: although the flat regime may converge faster at high accuracy, sufficiently large $L$ enhances feature separability, enabling accelerated training in the sharp regime.
@@
    Convergence Guarantee for ICL with Nonlinear Regression Functions: We provide formal convergence guarantees for a one-layer softmax attention model learning nonlinear regression functions. We prove that gradient descent achieves near-zero training loss in polynomial time across both flat and sharp $L$-regimes. We also characterize the two-phase training dynamics: an early phase in which attention scores between query tokens and target features rise rapidly, and a later phase in which these scores converge to one while attention to irrelevant features decays more slowly with oscillations. At convergence, query tokens consistently attend to highly relevant prompt tokens, enabling accurate predictions for previously unseen functions and demonstrating the ICL capability of transformers.
@@
    Novel Analysis Techniques: We develop new proof tools that explicitly connect the curvature of nonlinear task functions to the evolution of attention weights. In particular, we first decompose the prediction loss to explicitly relate it to attention weights and cross-feature gaps under nonlinear functions. We further show that the flat and sharp curvature regimes of the parameter $L$ lead to distinct gradient magnitudes, which in turn drive different convergence rates and shape the overall training dynamics. The impact of function curvature on the magnitude and stability of attention updates is, to our knowledge, not addressed in previous ICL literature [19, 27, 26].
@@ 2 Related Works
The Bayesian inference perspective, pioneered by [10] and further developed by [28, 29, 30], establishes a theoretical framework linking prompting strategies to probabilistic reasoning. This line of work interprets ICL as a form of implicit Bayesian model averaging, where transformers effectively perform approximated inference conditioned on the provided context. Another line of ICL work focuses on Markov chains to study the behavior of induction heads for transformers, which are designed to copy or compare tokens that follow previous occurrences [24, 31, 32, 33]. The central focus is to understand how transformers recover latent sequence structures [24] and transition rules [35, 36, 37], using token-level recurrence and dynamics.
@@
In contrast, the function learning perspective initiated by [11] demonstrated transformers’ remarkable ability to learn and interpolate simple function classes (particularly linear models) directly from context examples. This work sparked significant interest in understanding the mechanistic underpinnings of ICL, leading to important discoveries by [12] and [38], who revealed deep connections between attention mechanisms and gradient-based optimization dynamics. Recent advances have substantially expanded our understanding of ICL mechanisms: [39] provided a rigorous analysis of linear regression tasks, showing that trained transformers can implement both ridge regression gradient descent and exact least-squares solutions. [40] established comprehensive theoretical results encompassing expressive power, prediction capabilities, and sample complexity, while proposing general mechanisms for algorithmic selection. [26] interprets transformers as implementing functional gradient descent for nonlinear regression, but their analysis primarily focuses on the representational capacity and functional viewpoint of the learned predictor, without analyzing the ICL pre-training dynamics.
@@
Theoretical analysis of ICL learning dynamics. Recent theoretical work has made significant progress in understanding the pre-training dynamics of ICL in transformers, though important limitations remain [19, 20, 21, 22, 23, 41, 32, 42, 43, 16]. The foundational work by [19] established the first rigorous analysis of training dynamics for softmax attention in ICL settings, focusing on a single-head attention layer learning linear regression tasks. Their key theoretical result demonstrates that prompt tokens with features identical to the query token develop dominating attention weights during training. However, this analysis relies critically on strong assumptions about pairwise orthonormality of feature vectors, limiting its applicability to more general settings.
@@
Subsequent work extended these results to classification tasks [20], dual-head settings [42], and structured or hierarchical functions [22, 23]. Some studies explored the implicit bias of gradient descent and attention-based generalization [21, 41], while others examined information-theoretic and dynamical aspects of ICL [32, 43]. Despite these advances, most analyses still rely on restrictive assumptions, such as orthonormality features, fixed positional encoding, or carefully structured input distributions, limiting their ability to explain ICL under general tasks and practical learning conditions.
@@
Notations. In this paper, for a vector $\bm{v}$, we let $\|\bm{v}\|_{2}$ denote its $\ell$-22 norm. For some positive constant $C_{1}$ and $C_{2}$, we define $x=\Omega(y)$ if $x>C_{2}|y|$, $x=\Theta(y)$ if $C_{1}|y|<x<C_{2}|y|$, and $x=\mathcal{O}(y)$ if $x<c_{1}|y|$. We also denote by $x=o(y)$ if $x/y\rightarrow 0$. We use $\mathrm{poly}(C)$ to denote large constant degree polynomials of $C$. For a matrix $A$, we use $A_{i}$ to denote the $i$-th column of $A$, and $A_{i:j}$ to represent the collection of columns from the $i$-th to the $j$-th column (inclusive).
@@ 3.1 In-Context Learning Problem Setup
We consider a standard in-context learning (ICL) framework commonly used in prior studies [11, 19, 22]. The objective is to train a transformer model that can perform ICL over a designated class of functions $\mathcal{F}$, where each function $f\in\mathcal{F}$ corresponds to one task context. Here, we focus on nonlinear function classes further elaborated below.
Given each task (i.e., given one function $f$ randomly sampled from $\mathcal{F}$), a prompt with a sequence of $N$ input-response pairs $(x_{i},y_{i})$ as well as a query input $x_{\mathrm{query}}$ are sampled, where $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, and $y_{i}=f(x_{i})$. Let the input matrix $X=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}\\ \end{pmatrix}\in\mathbb{R}^{d\times N}$ and the response vector $\mathbf{y}=\begin{pmatrix}y_{1}&y_{2}&\cdots&y_{N}\end{pmatrix}\in\mathbb{R}^{1\times N}$. We adopt the following standard prompt embedding [11, 19, 22]:
$P=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathrm{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0\end{pmatrix}=\begin{pmatrix}X&x_{\mathrm{query}}\\ \mathbf{y}&0\end{pmatrix}\in\mathbb{R}^{(d+1)\times(N+1)}.$ (1)
@@
${\mathcal{F}=\left\{f:\;\begin{array}[]{l}|f(x)-f(x^{\prime})|\leq L\|x-x^{\prime}\|,\quad\forall\|x-x^{\prime}\|=\Theta(\delta_{0}),\\[6.99997pt] \forall v_{k}\in\mathbb{V},\exists\,v_{k^{\prime}}\in\mathbb{V},\;k^{\prime}\neq k,\;\text{such that }|f(v_{k})-f(v_{k^{\prime}})|=\Theta(L)\cdot\|v_{k}-v_{k^{\prime}}\|\end{array}\right\},}$ (4)
where $L>0$ and $\delta_{0}=\mathcal{O}(1)$. In particular, the functions in the class are said to satisfy non-degenerate $L$-Lipschitz condition, which imposes two natural requirements on the function class. First, the standard global $L$-Lipschitz condition ensures that the function $f$ does not change too rapidly, which is common in the literature and includes a wide class of linear and nonlinear functions. Second, the separation requirement guarantees that for any feature $v_{k}$ in the set $\mathbb{V}$, there exists at least one other feature $v_{k^{\prime}}$ such that the function difference between them achieves the order of variation defined by the Lipschitz constant. This ensures sufficient distinguishability of features by the function class for guaranteed learnability. This mild separation assumption is satisfied by almost any nontrivial $L$-Lipschitz function (e.g., scaled cosine, piecewise-polynomial, and small neural-network functions) and excludes only degenerate cases. Together, these two conditions ensure that the function class is sufficiently rich for ICL while avoiding unlearnable or trivial scenarios.
For each prompt $P$, the task-specific function $f(x)$ is independently drawn based on a task distribution $\mathcal{D}_{f}$, as long as $f(x)$ satisfies the property for the same $L$ and $\delta_{0}$ in Eq.˜4.
Feature Embeddings. Let $\mathbb{V}:=\big\{v_{k}\in\mathbb{R}^{d}\big|k=1,\cdots,K\big\}$ be the feature embeddings of tokens. For any $k\neq k^{\prime}$, we assume a separation of $\|v_{k}-v_{k^{\prime}}\|=\Theta(\Delta)$, where $\Delta=\Theta(1)$. Each data sample $x$ is modeled as a noisy perturbation of one of the vectors in $\mathbb{V}$. This assumption lets us control the separation $\Delta$ precisely, simplifying the analysis while retaining the essential geometry of the problem. Such a condition can be satisfied by various feature learning techniques to avoid feature collapse, e.g., disentangled representation learning [44, 45, 46]. We note that such a condition substantially generalizes the orthonormality assumption taken by the previous study [19, 20, 25, 24]. The prompt is sampled as follows. For a randomly chosen $v_{k}$, we assume $x$ satisfies $\|x-v_{k}\|=O(\epsilon_{x})$ with probability $p_{k}$, where $\epsilon_{x}=o(1)$ and $p_{k}=\Theta(\frac{1}{K})$. For analytical simplicity, we assume $x=v_{k}$ whenever this proximity condition holds. In our experiments in Section˜6, we further verify that our training dynamic analysis remains valid when tokens are drawn from general continuous distributions.